OpenStudy (anonymous):

Divide and simplify y/y^-5

OpenStudy (anonymous):

y^5/5 y^4

OpenStudy (anonymous):

\[y ^{5}/y = y ^{4}\]

OpenStudy (anonymous):

\[y*y^5=y^6\]

OpenStudy (anonymous):

A term with a negative expression is in inverse form. To simplify, invert the term, as I've shown, and solve

OpenStudy (anonymous):

\[y/y^-5\] \[y*y^5\] \[Y^6\]

OpenStudy (anonymous):

*negative exponent

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Disregard what I said. The other two were right, though they should have been clearer.\[y/y^{-5} = y/(1/y ^{5}) = y*y ^{5}/1 = y ^{6}\] You multiply numerators by the inverses of their denominators to evaluate such expressions